# Out of the box age estimation through facial imagery: A Comprehensive Benchmark of Vision-Language Models vs. out-of-the-box Traditional Architectures

### v1

@@ Abstract
Facial age estimation is critical for content moderation, age verification, and deepfake detection, yet no prior benchmark has systematically compared modern vision-language models (VLMs) against specialized age estimation architectures. We present the first large-scale cross-paradigm benchmark, evaluating 34 models—22 specialized architectures with publicly available pretrained weights and 12 general-purpose VLMs—across 8 standard datasets (UTKFace, IMDB-WIKI, MORPH, AFAD, CACD, FG-NET, APPA-REAL, AgeDB) totaling 1,100 test images per model. Our key finding is striking: zero-shot VLMs significantly outperform most specialized models, achieving an average MAE of 5.65 years compared to 9.88 for non-LLM models. The best VLM (Gemini 3 Flash Preview, MAE 4.32) outperforms the best non-LLM model (MiVOLO, MAE 5.10) by 15%. Only MiVOLO, which uniquely combines face and body features via Vision Transformers, competes with VLMs. We further analyze age verification at the 18-year threshold, revealing that non-LLM models exhibit 60–100% false adult rates on minors while VLMs achieve 13–25%, and demonstrate that coarse age binning (8–9 classes) consistently degrades MAE beyond 13 years. Our stratified analysis across 14 age groups reveals that all models struggle most at extreme ages (<<5 and 65+). These findings challenge the assumption that task-specific architectures are necessary for age estimation and suggest that the field should redirect toward distilling VLM capabilities into efficient specialized models.
@@ 1 Introduction
  • Age verification analysis: We evaluate all models at the critical 18-year threshold, revealing that most non-LLM models have unacceptably high false adult rates (60–100%) while VLMs achieve substantially lower rates (13–25%).
@@ 3.1 Models
VLMs are evaluated in a zero-shot setting via the OpenRouter API (accessed February 6–7, 2026) using a standardized prompt (Prompt B): “Estimate the age of the person in this photograph. Respond with ONLY a single integer representing their age in years. Do not include any other text, explanation, or units.” Temperature is set to 0 for deterministic outputs; maximum token limit is 256. Three additional prompt variants (A, C, D) were designed but not used for the main evaluation; see supplementary materials. We evaluate models from 9 providers: Google (Gemini), OpenAI (GPT-5), Anthropic (Claude), Alibaba (Qwen), Meta (Llama), Mistral, xAI (Grok), ByteDance (Seed), and Moonshot (Kimi). Model version strings are recorded in our code repository for reproducibility.
@@ 4.1 Overall Performance
Table 3 presents the full results across all 8 datasets (showing the top 17 and bottom 5; complete results in supplementary). The finding is unambiguous: zero-shot VLMs dominate. VLMs occupy 12 of the top 13 positions, achieving an average MAE of 5.65 years compared to 9.88 for non-LLM models—a 43% reduction in error.
@@
Figure 2 shows all 34 models ranked by average MAE. The performance gap between VLMs and non-LLM models is visually striking, with a clear separation around rank 13. Figure 3 further illustrates this distribution gap: the entire VLM interquartile range (5.21–6.02) falls below the non-LLM median (9.76).
@@
Figure 1 provides a detailed view of per-dataset MAE for all 34 models. Notable dataset-specific patterns include: (1) FG-NET exhibits the widest variance (2.36–28.11 MAE), as it contains very young children that many non-LLM models trained on adult-centric datasets fail to handle; (2) MORPH favors non-LLM models more than other datasets: MiVOLO achieves 4.02, likely due to MORPH’s controlled imaging conditions matching its training distribution; (3) CACD and IMDB-WIKI show the largest VLM advantage, suggesting that celebrity images with rich contextual cues (red carpet events, professional photos) particularly benefit VLMs’ holistic understanding; (4) AgeDB produces generally higher errors across all models (best: 5.32 MAE by Gemini 3 Flash), reflecting its challenging in-the-wild conditions and broad age range (1–101).
Table 3: Main results: MAE (years, $\downarrow$) across 8 datasets. Models sorted by average MAE. Orange = VLM, Blue = non-LLM. Bold = best per column; underline = second best. Only non-LLM models with publicly available pretrained weights are included; see supplementary for full details.
@@ 4.3 The Coarse Binning Problem
Four models in our benchmark use 8–9 discrete age categories (e.g., “0–2”, “4–6”, “8–12”, “15–20”, “25–32”, “38–43”, “48–53”, “60+”): FaceXFormer, ChienThan, AtulSingh, and Nixrajput. Table 4 shows that these models collectively achieve a mean MAE of 14.99, compared to 8.59 for CNN-regression models—a 75% increase in error attributable primarily to discretization.
@@
Category # Mean MAE Std.
VLM (zero-shot) 12 5.65 0.72
ViT (MiVOLO) 1 5.10
CNN-Regression 10 8.59 1.46
CNN-Expected Value 1 7.19
Multi-task 4 9.76 1.53
Retrieval (MWR) 1 10.04
Lightweight (SSR) 1 10.07
Coarse-bin (8–9 cls) 4 14.99 1.24
Non-LLM excl. coarse 18 8.74 1.80
@@ 4.4 Age Verification at the 18-Year Threshold
Figure 5 visualizes these error rates as a heatmap sorted by FAR. The pattern is clear: VLMs (top rows) show consistently low error rates across both columns, while non-LLM models (bottom rows) show high FAR values—a pattern directly attributable to their higher MAE around the threshold age range.
@@
Table 5: Age verification at the 18-year threshold across all datasets. FAR = False Adult Rate (% of actual minors predicted $\geq$18); FMR = False Minor Rate (% of actual adults predicted <<18). Lower FAR/FMR is better. Only models with $\geq$5 minor samples shown.
@@
Gemini 3 Flash Prev. 13.3 3.2 VLM
Gemini 2.5 Flash 14.7 2.8 VLM
GPT-5 Nano 16.0 3.6 VLM
Claude Sonnet 4.5 17.3 5.4 VLM
Qwen3-VL 235B 18.7 4.2 VLM
Seed-1.6 20.0 3.8 VLM
MiVOLO 21.3 4.1 Non-LLM
Kimi-K2.5 22.7 4.6 VLM
Mistral Small 3.2 24.0 5.1 VLM
DEX (VGG) 52.0 4.3 Non-LLM
SSR-Net 62.7 6.2 Non-LLM
DeepFace 65.3 5.9 Non-LLM
@@ 4.5 Age-Group Stratified Analysis
Figure 6 shows per-age-group MAE for the top 15 models. Several patterns emerge: (1) Children (<<5) and elderly (65+) are universally difficult, with even top VLMs producing MAE of 5–10 years in these ranges; (2) Young adults (20–34) are easiest for all models, with top VLMs achieving MAE <<3 years; (3) Middle-aged adults (35–64) show an interesting divergence: VLMs maintain MAE of 4–6 years while non-LLM models often exceed 10 years, suggesting VLMs leverage contextual cues (gray hair, wrinkles) more effectively.
@@ 5 Discussion
Our study has several limitations: (1) Sample size: we evaluate on 100 images per dataset (400 for AgeDB; 1,100 total per model), which provides reliable MAE estimates but may miss rare failure modes; (2) Single prompt: we define four prompt variants (A–D; see supplementary) but report results using only Prompt B. Prior work [24] shows VLM outputs can be sensitive to prompt phrasing; evaluating prompt sensitivity across all models is a natural extension; (3) No demographic bias analysis: we do not analyze performance disparities across gender, ethnicity, or skin tone, which is critical for fair deployment; (4) Snapshot evaluation: VLMs are updated frequently; our results reflect model versions available at evaluation time and may change with updates; (5) Label noise: some datasets (particularly IMDB-WIKI) have known label noise that affects all models equally but may inflate absolute MAE values; (6) Missing non-LLM baselines: recent specialized models trained on full datasets report substantially lower MAEs than our pretrained-model evaluation (e.g., MSDNN achieves 2.59 on MORPH-II [4]; Paplham and Franc [33] report comparable results with optimized splits). Our benchmark restricts non-LLM models to those with publicly available pretrained weights and evaluates them on a fixed random sample rather than models retrained on full datasets, so direct comparison of our sampled MAEs to published full-dataset results is not appropriate. Including additional recent architectures with public weights would strengthen future versions of this benchmark.

### v2

@@ Abstract
Facial age estimation is critical for content moderation, age verification, and deepfake detection, yet no prior benchmark has systematically compared modern vision-language models (VLMs) against specialized age estimation architectures. We present the first large-scale cross-paradigm benchmark, evaluating 34 models—22 specialized architectures with publicly available pretrained weights and 12 general-purpose VLMs—across 8 standard datasets (UTKFace, IMDB-WIKI, MORPH, AFAD, CACD, FG-NET, APPA-REAL, AgeDB) totaling 1,100 test images per model. Our key finding is striking: zero-shot VLMs significantly outperform most specialized models, achieving an average MAE of 5.65 years compared to 9.88 for non-LLM models. The best VLM (Gemini 3 Flash Preview, MAE 4.32) outperforms the best non-LLM model (MiVOLO, MAE 5.10) by 15%. Only MiVOLO, which uniquely combines face and body features via Vision Transformers, competes with VLMs. We further analyze age verification at the 18-year threshold, revealing that most non-LLM models exhibit 39–100% false adult rates on minors while VLMs achieve 16–29%, and demonstrate that coarse age binning (8–9 classes) consistently degrades MAE beyond 13 years. Our stratified analysis across 14 age groups reveals that all models struggle most at extreme ages (<<5 and 65+). These findings challenge the assumption that task-specific architectures are necessary for age estimation and suggest that the field should redirect toward distilling VLM capabilities into efficient specialized models.
@@ 1 Introduction
  • Age verification analysis: We evaluate all models at the critical 18-year threshold, revealing that most non-LLM models have unacceptably high false adult rates (39–100%) while VLMs achieve substantially lower rates (16–29%).
@@ 3.1 Models
VLMs are evaluated in a zero-shot setting via the OpenRouter API (accessed February 6–7, 2026) using a standardized prompt: “Estimate the age of the person in this photograph. Respond with ONLY a single integer representing their age in years. Do not include any other text, explanation, or units.” Temperature is set to 0 for deterministic outputs; maximum token limit is 256. We evaluate models from 9 providers: Google (Gemini), OpenAI (GPT-5), Anthropic (Claude), Alibaba (Qwen), Meta (Llama), Mistral, xAI (Grok), ByteDance (Seed), and Moonshot (Kimi). Model version strings are recorded in our code repository for reproducibility.
@@ 4.1 Overall Performance
Table 3 presents the full results across all 8 datasets (showing the top 17 and bottom 5). The finding is unambiguous: zero-shot VLMs dominate. VLMs occupy 12 of the top 13 positions, achieving an average MAE of 5.65 years compared to 9.88 for non-LLM models—a 43% reduction in error.
@@
Figure 1 shows all 34 models ranked by average MAE. The performance gap between VLMs and non-LLM models is visually striking, with a clear separation around rank 13. Figure 2 further illustrates this distribution gap: the entire VLM interquartile range (5.21–6.06) falls below the non-LLM median (9.45).
@@
Figure 3 provides a detailed view of per-dataset MAE for all 34 models. Notable dataset-specific patterns include: (1) FG-NET exhibits the widest variance (2.36–20.94 MAE), as it contains very young children that many non-LLM models trained on adult-centric datasets fail to handle; (2) MORPH favors non-LLM models more than other datasets: MiVOLO achieves 4.02, likely due to MORPH’s controlled imaging conditions matching its training distribution; (3) CACD and IMDB-WIKI show the largest VLM advantage, suggesting that celebrity images with rich contextual cues (red carpet events, professional photos) particularly benefit VLMs’ holistic understanding; (4) AgeDB produces generally higher errors across all models (best: 5.32 MAE by Gemini 3 Flash), reflecting its challenging in-the-wild conditions and broad age range (1–101).
Table 3: Main results: MAE (years, $\downarrow$) across 8 datasets. Models sorted by average MAE. Orange = VLM, Blue = non-LLM. Bold = best per column; underline = second best. Only non-LLM models with publicly available pretrained weights are included.
@@
Category # Mean MAE Std.
VLM (zero-shot) 12 5.65 0.72
ViT (MiVOLO) 1 5.10
CNN-Regression 10 8.52 1.36
CNN-Expected Value 1 7.19
Multi-task 4 9.94 1.62
Retrieval (MWR) 1 10.04
Lightweight (SSR) 1 10.07
Coarse-bin (8–9 cls) 4 15.00 1.25
Non-LLM excl. coarse 18 8.74 1.71
@@ 4.3 The Coarse Binning Problem
Four models in our benchmark use 8–9 discrete age categories (e.g., “0–2”, “4–6”, “8–12”, “15–20”, “25–32”, “38–43”, “48–53”, “60+”): FaceXFormer, ChienThan, AtulSingh, and Nixrajput. Table 4 shows that these models collectively achieve a mean MAE of 15.00, compared to 8.52 for CNN-regression models—a 76% increase in error attributable primarily to discretization.
@@ 4.4 Age Verification at the 18-Year Threshold
Figure 5 visualizes these error rates as a heatmap sorted by FAR. A notable outlier is FaceXFormer, which achieves the lowest FAR (1.9%) but at the cost of an extreme FMR (49.8%)—it effectively predicts nearly all individuals as minors, rendering it useless for practical age verification. Excluding this degenerate case, the pattern is clear: VLMs cluster near the top with consistently low error rates across both columns, while non-LLM models dominate the bottom rows with high FAR values—a pattern directly attributable to their higher MAE around the threshold age range.
@@
Table 5: Age verification at the 18-year threshold across all datasets. FAR = False Adult Rate (% of actual minors predicted $\geq$18); FMR = False Minor Rate (% of actual adults predicted <<18). Lower FAR/FMR is better. Selected models with $\geq$5 minor samples shown.
@@
Seed-1.6 16.3 3.5 VLM
Gemini 3 Flash Prev. 19.2 1.4 VLM
Qwen3-VL 235B 19.2 2.7 VLM
GPT-5 Nano 20.4 1.7 VLM
Mistral Small 3.2 21.2 2.6 VLM
MiVOLO 21.2 2.5 Non-LLM
Gemini 2.5 Flash 22.1 1.8 VLM
Claude Sonnet 4.5 25.0 2.9 VLM
Kimi-K2.5 28.8 2.1 VLM
Herosan-Age 38.5 3.1 Non-LLM
DEX (VGG) 60.6 1.2 Non-LLM
@@ 4.5 Age-Group Stratified Analysis
Figure 6 shows per-age-group MAE for the top 15 models, revealing distinct performance patterns across the human lifespan.
(1) Extreme Age Groups (<<5 and 65+): These ranges are universally difficult, with even top-tier VLMs producing MAE of 5–10 years. This performance degradation likely stems from a long-tail data distribution issue: in typical internet-scale pretraining sets and social media, images of infants and the elderly are significantly underrepresented compared to young adults.
(2) Young Adults (20–34): This cohort is the “sweet spot” for all models, with top VLMs achieving MAE <<3 years. We hypothesize this is due to massive exposure in training data, as this age group is the most active on digital platforms and social media. Furthermore, individuals in this range rarely undergo significant age-altering cosmetic procedures, preserving natural aging cues.
(3) Middle-aged Adults (35–64) and the VLM Divergence: An interesting divergence emerges here: VLMs maintain a stable MAE of 4–6 years, while non-LLM models often exceed 10 years. We propose two complementary hypotheses for this VLM superiority:
  • Contextual Reasoning: VLMs can leverage non-facial cues (e.g., gray hair, subtle wrinkles, professional attire, or “red carpet” backgrounds in celebrity datasets like IMDB-WIKI) to infer age holistically.
  • Robustness to Cosmetic Alterations: This age group is more likely to undergo cosmetic surgery or use heavy anti-aging treatments. While such alterations may confuse non-LLM models that rely strictly on low-level facial textures, VLMs may “cheat” by associating the person’s identity with known biographical timelines or by identifying the specific visual style of different eras, effectively bypassing the deceptive effects of cosmetic procedures.
@@ 5 Discussion
Our study has several limitations: (1) Sample size: we evaluate on 100 images per dataset (400 for AgeDB; 1,100 total per model), which provides reliable MAE estimates but may miss rare failure modes; (2) Single prompt: results are reported using a single prompt. Prior work [24] shows VLM outputs can be sensitive to prompt phrasing; evaluating prompt sensitivity across all models is a natural extension; (3) No demographic bias analysis: we do not analyze performance disparities across gender, ethnicity, or skin tone, which is critical for fair deployment; (4) Snapshot evaluation: VLMs are updated frequently; our results reflect model versions available at evaluation time and may change with updates; (5) Label noise: some datasets (particularly IMDB-WIKI) have known label noise that affects all models equally but may inflate absolute MAE values; (6) Missing non-LLM baselines: recent specialized models trained on full datasets report substantially lower MAEs than our pretrained-model evaluation (e.g., MSDNN achieves 2.59 on MORPH-II [4]; Paplham and Franc [33] report comparable results with optimized splits). Our benchmark restricts non-LLM models to those with publicly available pretrained weights and evaluates them on a fixed random sample rather than models retrained on full datasets, so direct comparison of our sampled MAEs to published full-dataset results is not appropriate. Including additional recent architectures with public weights would strengthen future versions of this benchmark.